# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 3.1 Cross-fit procedure
The next subgaussian condition is unnecessary for obtaining a Berry-Esseen-type bound, but it may be reasonable and leads to tighter bounds. For any function $Q$ and $a^{\prime}\in\{0,1\}$, define function $H_{a^{\prime}}(Q):v=(x,a,y)\mapsto{\mathcal{T}}_{a^{\prime}}(Q)(v)-{\mathcal{T}}_{a^{\prime}}(Q_{\#,a^{\prime}})(v)=\{1-{\mathbbm{1}}(a=a^{\prime})\}\{Q(x)-Q_{\#,a^{\prime}}(x)\}/\pi_{*}(a^{\prime}\mid x)$.
Condition 7 (Subgaussian $\hat{Q}_{k,a}$).
For all $k\in\{1,\ldots,K\}$ and $a\in\{0,1\}$, the followings hold.
  1. 1.
    Let $V=(X,A,Y)$ denote a random vector that is independent of the data $(V_{i})_{i=1}^{n}$ and identically distributed as a data point $V_{i}$. For $P_{*}$-almost all $\hat{Q}_{k,a}$, $H_{a}(\hat{Q}_{k,a})(V)/\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}\mid\hat{Q}_{k,a}$ is subgaussian with a variance proxy $\hat{s}_{1}^{2}$ (namely a subgaussian parameter $\hat{s}_{1}>0$).
  2. 2.
    $(\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}-{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2})/\sqrt{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}$ is subgaussian with a variance proxy $\hat{s}_{2}^{2}$ (namely a subgaussian parameter $\hat{s}_{2}>0$).
Subgaussian conditions are common in high-dimensional statistics. An overview of subgaussian distributions can be found in, for example, Wainwright, (2019). To the best of my knowledge, subgaussian properties of flexible machine-learning-based outcome model estimators have rarely been established, despite established convergence rates (e.g., Chen and White,, 1999; Bühlmann and Yu,, 2003; Benkeser and Van Der Laan,, 2016). Still, Condition 7 might be reasonable. As shown in Lemma S5 in the Supplementary Material, $H_{a}(Q)(V)$ has mean zero and variance of order $\|Q-Q_{\#,a}\|_{2}^{2}$. Thus, Part 1 may hold if, for $P_{*}$-almost all $\hat{Q}_{k,a}$, $\hat{Q}_{k,a}(X)-Q_{\#,a}(X)\mid\hat{Q}_{k,a}$ has a light tail. In Part 2, the denominator $\sqrt{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}$ is an upper bound of the standard deviation of the random variable $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}$ in the numerator, and this part may hold if $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}$ has a light tail. Another similar subgaussian condition S1 on $P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}$ is listed in Section S1 of the Supplementary Material.
For readability, I next present an asymptotic Berry-Esseen-type bound for the cross-fit Wald-CI in the scenario $n\to\infty$, with $P_{*}$ and all constants in Conditions 15, 7, and S1 fixed. A more intricate non-asymptotic version is presented in Theorem S1 in Section S2 of the Supplementary Material.
@@
$\left|\Pr(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$={\mathrm{O}}\left(K^{2/3}\left\{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3}+\sqrt{\frac{K\log n}{n}}\right).$ (1)
Additionally under subgaussian conditions 7 and S1, (1) holds with its right-hand side replaced by
${\mathrm{O}}\left(\sqrt{K{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\log n+\sqrt{\frac{K\log n}{n}}\right).$
The bound for a $(1-\alpha)$-level two-sided Wald-CI can be derived by noting that $\Pr(-z_{\alpha/2}\hat{\sigma}_{a}\leq\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha/2}\hat{\sigma}_{a})=\Pr(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha/2}\hat{\sigma}_{a})-\Pr(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})<-z_{\alpha/2}\hat{\sigma}_{a})$. The proof of of this theorem can be found in Section S4.2 of the Supplementary Material. It is based on a more general version of the delta method for CI coverage (Section 2.7 in Hall,, 1992), a Taylor’s expansion to analyze the effect of the bias $\sigma_{\dagger,a}-\sigma_{\#,a}$ of the variance estimator, non-asymptotic bounds on the deviation of $\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})$ from its linear approximation $\sqrt{n}(P_{n}-P_{*})D_{a}(Q_{\#,a},\psi_{*,a})$, the Berry-Esseen Theorem to bound the deviation of the linear approximation $\sqrt{n}(P_{n}-P_{*})D_{a}(Q_{\#,a},\psi_{*,a})$ from the asymptotic normal distribution, and a non-asymptotic tail bound of $\hat{\sigma}_{a}-\sigma_{\dagger,a}$. Because I use the crude union bound to account for cross-fitting, the rate might be improved regarding $K$.
@@
Condition 8 (Consistency or common mean for $\hat{Q}_{k,a}$).
@@
    For all $k=1,\ldots,K$, $Q_{\#,a}(X)={\mathbb{E}}_{\hat{Q}_{k,a}}[\hat{Q}_{k,a}(X)]$ $P_{*}$-almost surely.
Part 1 of Condition 8 holds if $\hat{Q}_{k,a}$ is consistent for the true outcome model $Q_{*,a}$. Part 2 of Condition 8 holds if Condition 6 holds and $Q_{\#,a}$ is taken to be the common mean of all outcome model estimators $\hat{Q}_{k,a}$, regardless of whether the outcome model estimator $\hat{Q}_{k,a}$ converges to the truth $Q_{*,a}$.
@@
Under Conditions 1 and 3, as $n\to\infty$, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}(\sqrt{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}+n^{-1})$. Additionally under Condition 8, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}({\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+n^{-1})$.
A non-asymptotic version is presented in Theorem S2 in Section S2 of the Supplementary Material. The term involving ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ arises from explicitly calculating or bounding the variance of ${\mathcal{T}}_{a}(\hat{Q}_{k,a})$, which contributes to $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}$. Under Condition 8, the rate is faster because, when the propensity score $\pi_{*}$ is known, $Q_{\#,a}$ is a projection of $\hat{Q}_{k,a}$ or $Y$, so that the linear contribution of $\hat{Q}_{k,a}-Q_{\#,a}$ to $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}$ vanishes, and thus $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}$ is almost quadratic. A formal proof can be found in Section S4.2 of the Supplementary Material. This theorem implies that, regardless of whether Condition 8 holds, the term involving $\sigma_{\dagger,a}-\sigma_{\#,a}$ in (1) is not a leading term in the Berry-Esseen-type bound in Theorem 1. I defer further interpretation of Theorem 1 to Section 3.3.
@@ 3.2 Non-cross-fit procedure
A Donsker condition is needed for $\tilde{\psi}_{a}$ to be asymptotically normal, but slightly stronger conditions appear necessary to obtain more insights like convergence rates into CI coverage than a mere convergence of CI coverage to the nominal coverage. For concreteness and illustration, I consider the following sufficient condition of a Donsker condition based on covering number and uniform entropy integral of a uniformly bounded function class (van der Vaart and Wellner,, 1996). For any function class $\mathcal{G}$}, let $N(\epsilon,{\mathcal{G}},\|\cdot\|)$ denote the $\epsilon$-covering number of $\mathcal{G}$} with respect to a pseudo-norm $\|\cdot\|$ (the minimal number of balls with radius $\epsilon$ that altogether cover $\mathcal{G}$}), and
$J(\delta,{\mathcal{G}},G):=\int_{0}^{\delta}\sup_{\mu}\sqrt{1+\log N(\epsilon\|G\|_{2,\mu},{\mathcal{G}},L_{2}(\mu))}{\mathrm{d}}\epsilon$
@@
Condition 9 (Sufficient conditions for Donsker condition).
@@
In the asymptotic scenario, I assume that $\mathcal{F}$} is fixed and does not depend on $n$. I also consider the following subgaussian condition similar to Condition 7, and another subgaussian condition S2 in Section S1 of the Supplementary Material that is similar to Condition S1.
Condition 10 (Subgaussian $\tilde{Q}_{a}$).
$(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})/\sqrt{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}$ is subgaussian with a variance proxy $\tilde{s}_{1}^{2}$ (namely a subgaussian parameter $\tilde{s}_{1}^{2}$).
@@
Under Conditions 15 and 9, for any $\delta>0$ that may depend on $n$,
$\left|\Pr(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$={\mathrm{O}}\left(\tilde{\Delta}(\delta)+J(2\delta,{\mathcal{F}},M)+\frac{J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}\sqrt{n}}+(\delta+n^{-1/2})\log n\right),$ (2)
where $\tilde{\Delta}(\delta):={\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}/\delta^{2}+({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/3}$. Additionally under subgaussian conditions 10 and S2, if $2M\delta>{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$, then (2) holds with $\tilde{\Delta}(\delta)$ replaced by
$\tilde{\Delta}^{\prime}(\delta):=\exp\left\{-\frac{(2\delta M-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}{2\tilde{s}_{1}^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}+\sqrt{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n}.$
Although Conditions 2 and 5 are implied by the outcome boundedness in Condition 9, I include these two redundant conditions in Theorem 3 to make the conditions appear more similar to those in Theorem 1. Since $J(2\delta,{\mathcal{F}},M)\geq 2\delta$, the rate’s upper bound in the right-hand side of (2) is lower bounded by ${\mathrm{O}}(\{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\}^{1/3}\{\log n\}^{2/3}+n^{-1/2}\log n)$ without subgaussian conditions, and lower bounded by ${\mathrm{O}}(\sqrt{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n}+n^{-1/2}\log n)$ (up to log factors) with subgaussian conditions, which are comparable to Theorem 1 with fixed $K$. Theorem S3 can be shown with an argument similar to Theorem S1, except that the additional empirical process terms are bounded by Theorems 5.1 and 5.2 in Chernozhukov et al., (2014).
I next show a similar result on $|\sigma_{\dagger,a}-\sigma_{\#,a}|$ to Theorem 2, with a non-asymptotic version, Theorem S4 in Section S2 of the Supplementary Material. I consider a similar condition to Condition 8.
Condition 11 (Consistency of $\tilde{Q}_{a}$ or suitable $Q_{\#,a}$).
@@
    $Q_{\#,a}(X)={\mathbb{E}}_{\tilde{Q}_{a}}[\tilde{Q}_{a}(X)]$ $P_{*}$-almost surely.
Part 1 of Condition 11 holds if $\tilde{Q}_{a}$ is consistent for $Q_{*,a}$, similarly to Part 1 of Condition 8. Part 2 of Condition 11 is even weaker than Part 2 of Condition 8 as it always holds for an appropriately chosen $Q_{\#,a}$.
@@
Under Conditions 1, 3, 4 and 9, as $n\to\infty$, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}(\sqrt{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}+n^{-1/2})$. Additionally under Condition 11, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}+n^{-1/2})$.
It is possible to show that the ${\mathrm{O}}(n^{-1/2})$ term is ${\mathrm{o}}(n^{-1/2})$, with the exact rate depending on the covering number of $\mathcal{F}$}. This result shows that the term involving $\sigma_{\dagger,a}-\sigma_{\#,a}$ in (2) does not dominate in the Berry-Esseen-type bound in Theorem 3. If one is willing to ignore log factors when interpreting rates and ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}={\mathrm{O}}(n^{-1})$ (e.g., $\tilde{Q}_{a}$ is estimated in a parametric model), then $\sigma_{\dagger,a}-\sigma_{\#,a}={\mathrm{O}}(n^{-1/2})$ is at most comparable to $\sqrt{\log n/n}$ in (2).
In many cases, the bounded uniform entropy integral condition $J(1,{\mathcal{F}},M)<\infty$ in Condition 9 is verified by establishing that $\mathcal{F}$} is subset of a VC or VC-hull class. I adopt the definitions in van der Vaart and Wellner, (1996), where a VC index is an integer greater than one, and consider the following VC or VC-hull condition.
Condition 12 (VC or VC-hull).
@@
Under Part 1 of Condition 12, $\sup_{\mu}\log N(\epsilon M,{\mathcal{F}},L_{2}(\mu))\leq C_{1}(\nu)+2(\nu-1)\log(1/\epsilon)$ for all $\epsilon\in(0,1)$, where $C_{1}(\nu)$ is a constant depending on $\nu$ only (Theorem 2.6.7 in van der Vaart and Wellner,, 1996). Under Part 2 of Condition 12, $\sup_{\mu}\log N(\epsilon M,{\mathcal{F}},L_{2}(\mu))\leq C_{2}(\nu)\epsilon^{-2+2/\nu}$ for all $\epsilon\in(0,1)$, where $C_{2}(\nu)$ is another constant depending on $\nu$ only (Corollary 2.6.12 in van der Vaart and Wellner,, 1996). These lead to the following result on the rates for (2), depending on which part of Condition 12 holds, and whether Conditions 10 and S2 hold.
@@
Suppose that Conditions 15 and 9 hold. Consider the asymptotic scenario $n\to\infty$.

### v2

@@ 3.1 Cross-fit procedure
The next approximate sub-Weibull conditions are unnecessary for obtaining a Berry-Esseen-type bound, but it may be reasonable and leads to tighter bounds.
Condition 7 (Approximate sub-Weibull tail of $\hat{Q}_{k,a}$).
For all $k\in\{1,\ldots,K\}$ and $a\in\{0,1\}$, there exist constants $\hat{a}_{1},\hat{b}_{1},\hat{d}_{1}\geq 0,\hat{c}_{1},\hat{q}_{1}>0$ and a deterministic positive function $\hat{r}_{1}$ such that, for all $t>0$,
${\mathrm{pr}}\left(\frac{\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}}{\hat{r}_{1}(n)}>t\right)\leq\hat{a}_{1}n^{\hat{d}_{1}}\exp\left(-\hat{c}_{1}t^{\hat{q}_{1}}\right)+\hat{b}_{1}(\log n/n)^{1/2}.$
Condition 7 requires that the $L_{2}(P_{*})$-distance between $\hat{Q}_{k,a}$ and $Q_{\#,a}$, normalized by a rate $\hat{r}_{1}(n)$, is roughly sub-Weibull, with a possible multiplicative factor that grows at a polynomial rate at most and a possible relaxation term that vanishes faster than $n^{-1/2}$. The rate $\hat{r}_{1}(n)$ can be (an upper bound of) the convergence rate of $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}$. The $(\log n/n)^{1/2}$ term allows for any rare event whose probability decays with $n$ rapidly, for example, the event that all individuals are in the other treatment arm so that there is no data to obtain $\hat{Q}_{k,a}$. If one wishes to control the tail of $\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}$ via asymptotic normality, this $(\log n/n)^{1/2}$ term can also handle normal approximation errors to be controlled by a Berry-Esseen-type bound. I illustrate a concrete example of an outcome model estimator satisfying this sub-Weibull tail bound, series regression, in Section S2 of the Supplementary Material. For empirical risk minimizers, such approximate sub-Weibull tail bounds might be obtained via empirical process tail bounds (e.g., Theorems 2.14.36, 2.14.37, 2.15.5, 2.15.11 in van der Vaart and Wellner,, 2023).
Condition 8 (Approximate sub-Weibull tail of $P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}$).
There exist constants $\hat{a}_{2},\hat{b}_{2},\hat{d}_{2}\geq 0,\hat{c}_{2},\hat{q}_{2}>0$ and a deterministic positive function $\hat{r}_{2}$ such that, for all $k\in\{1,\ldots,K\}$, $a\in\{0,1\}$, and $t>0$,
${\mathrm{pr}}\left(\frac{|P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{E}[P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]|}{\hat{r}_{2}(n)}>t\right)\leq\hat{a}_{2}n^{\hat{d}_{2}}\exp\left(-\hat{c}_{2}t^{\hat{q}_{2}}\right)+\hat{b}_{2}(\log n/n)^{1/2}.$
As shown in Lemma S6 in the Supplementary Material, because of the integration with respect to $P_{*}$ and the boundedness of $\hat{Q}_{k,a}$ in Condition 4, the random variable in the numerator has a tail similar to the $L_{2}(P_{*})$-distance between $\hat{Q}_{k,a}$ and its mean. Thus, if $Q_{\#,a}$ is the mean of $\hat{Q}_{k,a}$ and Condition 7 holds, Condition 8 may hold with $\hat{r}_{2}=\hat{r}_{1}$.
For readability, I next present an asymptotic Berry-Esseen-type bound for the cross-fit Wald-CI in the scenario $n\to\infty$, with $P_{*}$ and all constants in Conditions 15, 7, and 8 fixed. A more intricate non-asymptotic version is presented in Theorem S1 in Section S1 of the Supplementary Material.
@@
$\left|{\mathrm{pr}}(n^{1/2}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$={\mathrm{O}}\left(K^{2/3}\left\{{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3}+\left\{\frac{K\log n}{n}\right\}^{1/2}\right).$ (1)
Additionally under approximate sub-Weibull conditions 7 and 8, (1) holds with its right-hand side replaced by
${\mathrm{O}}\left(K^{1/2}\hat{r}_{1}(n)(\log n)^{(\hat{q}_{1}+2)/(2\hat{q}_{1})}+\hat{r}_{2}(n)(\log n)^{1/\hat{q}_{2}}+K\{\log n/n\}^{1/2}\right).$
The bound for a $(1-\alpha)$-level two-sided Wald-CI can be derived by noting that ${\mathrm{pr}}(-z_{\alpha/2}\hat{\sigma}_{a}\leq n^{1/2}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha/2}\hat{\sigma}_{a})={\mathrm{pr}}(n^{1/2}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha/2}\hat{\sigma}_{a})-{\mathrm{pr}}(n^{1/2}(\hat{\psi}_{a}-\psi_{*,a})<-z_{\alpha/2}\hat{\sigma}_{a})$. The proof of of this theorem can be found in Section S5.2 of the Supplementary Material. It is based on a more general version of the delta method for CI coverage (Section 2.7 in Hall,, 1992), a Taylor’s expansion to analyze the effect of the bias $\sigma_{\dagger,a}-\sigma_{\#,a}$ of the variance estimator, non-asymptotic bounds on the deviation of $n^{1/2}(\hat{\psi}_{a}-\psi_{*,a})$ from its linear approximation $n^{1/2}(P_{n}-P_{*})D_{a}(Q_{\#,a},\psi_{*,a})$, the Berry-Esseen Theorem to bound the deviation of the linear approximation $n^{1/2}(P_{n}-P_{*})D_{a}(Q_{\#,a},\psi_{*,a})$ from the asymptotic normal distribution, and a non-asymptotic tail bound of $\hat{\sigma}_{a}-\sigma_{\dagger,a}$. Because I use the crude union bound to account for cross-fitting, the rate might be improved regarding $K$.
@@
Condition 9 (Consistency or common mean for $\hat{Q}_{k,a}$).
@@
    For all $k=1,\ldots,K$, $Q_{\#,a}(X)={E}_{\hat{Q}_{k,a}}[\hat{Q}_{k,a}(X)]$ $P_{*}$-almost surely.
Part 1 of Condition 9 holds if $\hat{Q}_{k,a}$ is consistent for the true outcome model $Q_{*,a}$. Part 2 of Condition 9 holds if Condition 6 holds and $Q_{\#,a}$ is taken to be the common mean of all outcome model estimators $\hat{Q}_{k,a}$, regardless of whether the outcome model estimator $\hat{Q}_{k,a}$ converges to the truth $Q_{*,a}$.
@@
Under Conditions 1 and 3, as $n\to\infty$, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}(\{{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\}^{1/2}+n^{-1})$. Additionally under Condition 9, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}({E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+n^{-1})$.
A non-asymptotic version is presented in Theorem S2 in Section S1 of the Supplementary Material. The term involving ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ arises from explicitly calculating or bounding the variance of ${\mathcal{T}}_{a}(\hat{Q}_{k,a})$, which contributes to $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}$. Under Condition 9, the rate is faster because, when the propensity score $\pi_{*}$ is known, $Q_{\#,a}$ is a projection of $\hat{Q}_{k,a}$ or $Y$, so that the linear contribution of $\hat{Q}_{k,a}-Q_{\#,a}$ to $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}$ vanishes, and thus $\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}$ is almost quadratic. A formal proof can be found in Section S5.2 of the Supplementary Material. This theorem implies that, regardless of whether Condition 9 holds, the term involving $\sigma_{\dagger,a}-\sigma_{\#,a}$ in (1) is not a leading term in the Berry-Esseen-type bound in Theorem 1. I defer further interpretation of Theorem 1 to Section 3.3.
@@ 3.2 Non-cross-fit procedure
A Donsker condition is needed for $\tilde{\psi}_{a}$ to be asymptotically normal, but slightly stronger conditions appear necessary to obtain more insights like convergence rates into CI coverage than a mere convergence of CI coverage to the nominal coverage. For concreteness and illustration, I consider the following sufficient condition of a Donsker condition based on covering number and uniform entropy integral of a uniformly bounded function class (van der Vaart and Wellner,, 2023). For any function class $\mathcal{G}$}, let $N(\epsilon,{\mathcal{G}},\|\cdot\|)$ denote the $\epsilon$-covering number of $\mathcal{G}$} with respect to a pseudo-norm $\|\cdot\|$ (the minimal number of balls with radius $\epsilon$ that altogether cover $\mathcal{G}$}), and
$J(\delta,{\mathcal{G}},G):=\int_{0}^{\delta}\sup_{\mu}\left\{1+\log N\big(\epsilon\|G\|_{2,\mu},{\mathcal{G}},L_{2}(\mu)\big)\right\}^{1/2}{\mathrm{d}}\epsilon$
@@
Condition 10 (Sufficient conditions for Donsker condition).
@@
In the asymptotic scenario, I assume that $\mathcal{F}$} is fixed and does not depend on $n$. I also consider the following approximate sub-Weibull conditions similar to Conditions 7 and 8.
Condition 11 (Approximate sub-Weibull tail of $\tilde{Q}_{a}$).
There exist constants $\tilde{a}_{1},\tilde{b}_{1},\tilde{d}_{1}\geq 0,\tilde{c}_{1},\tilde{q}_{1}>0$ and a deterministic positive function $\tilde{r}_{1}$ such that, for all $t>0$,
${\mathrm{pr}}\left(\frac{\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}}{\tilde{r}_{1}(n)}>t\right)\leq\tilde{a}_{1}n^{\tilde{d}_{1}}\exp\left(\tilde{c}_{1}t^{\tilde{q}_{1}}\right)+\tilde{b}_{1}n^{-1/2}\log n.$
Condition 12 (Approximate sub-Weibull tail of $P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}$).
There exist constants $\tilde{a}_{2},\tilde{b}_{2},\tilde{d}_{2}\geq 0,\tilde{c}_{2},\tilde{q}_{2}>0$ and a deterministic positive function $\tilde{r}_{2}$ such that, for all $a\in\{0,1\}$ and $t>0$,
${\mathrm{pr}}\left(\frac{|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{E}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|}{\tilde{r}_{2}(n)}>t\right)\leq\tilde{a}_{2}n^{\tilde{d}_{2}}\exp\left(-\tilde{c}_{2}t^{\tilde{q}_{2}}\right)+\tilde{b}_{2}n^{-1/2}\log n.$
@@
Under Conditions 15 and 10, for any $\delta>0$ that may depend on $n$,
$\left|{\mathrm{pr}}(n^{1/2}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$={\mathrm{O}}\left(\tilde{\Delta}(\delta)+J(2\delta,{\mathcal{F}},M)+\frac{J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}n^{1/2}}+(\delta+n^{-1/2})\log n\right),$ (2)
where $\tilde{\Delta}(\delta):={E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}/\delta^{2}+({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/3}$. Additionally under approximate sub-Weibull conditions 11 and 12, (2) holds with $\tilde{\Delta}(\delta)$ replaced by
$\tilde{\Delta}^{\prime}(\delta):=n^{\tilde{d}_{1}}\exp\left\{-\tilde{c}_{1}\left(\frac{2\delta M}{\tilde{r}_{1}(n)}\right)^{\tilde{q}_{1}}\right\}+\tilde{r}_{2}(n)(\log n)^{1/\tilde{q}_{2}}.$
The rates in Theorem 3 might not be intuitive, so I will show the rates for VC-subgraph and VC-hull classes in Corollary 1 later in this section. Although Conditions 2 and 5 are implied by the outcome boundedness in Condition 10, I include these two redundant conditions in Theorem 3 to make the conditions appear more similar to those in Theorem 1. Since $J(2\delta,{\mathcal{F}},M)\geq 2\delta$, the rate’s upper bound in the right-hand side of (2) is lower bounded by ${\mathrm{O}}(\{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\}^{1/3}+n^{-1/2})$ (up to log factors) without approximate sub-Weibull conditions, and lower bounded by ${\mathrm{O}}(\{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\}^{1/2}+n^{-1/2})$ (up to log factors) with approximate sub-Weibull conditions, which are comparable to Theorem 1 with fixed $K$. Theorem S3 can be shown with an argument similar to Theorem S1, except that the additional empirical process terms are bounded by Theorems 5.1 and 5.2 in Chernozhukov et al., (2014).
I next show a similar result on $|\sigma_{\dagger,a}-\sigma_{\#,a}|$ to Theorem 2, with a non-asymptotic version, Theorem S4 in Section S1 of the Supplementary Material. I consider a similar condition to Condition 9.
Condition 13 (Consistency of $\tilde{Q}_{a}$ or suitable $Q_{\#,a}$).
@@
    $Q_{\#,a}(X)={E}_{\tilde{Q}_{a}}[\tilde{Q}_{a}(X)]$ $P_{*}$-almost surely.
Part 1 of Condition 13 holds if $\tilde{Q}_{a}$ is consistent for $Q_{*,a}$, similarly to Part 1 of Condition 9. Part 2 of Condition 13 is even weaker than Part 2 of Condition 9 as it always holds for an appropriately chosen $Q_{\#,a}$.
@@
Under Conditions 1, 3, 4 and 10, as $n\to\infty$, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}(\{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\}^{1/2}+n^{-1/2})$. Additionally under Condition 13, $|\sigma_{\dagger,a}-\sigma_{\#,a}|={\mathrm{O}}({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}+n^{-1/2})$.
It is possible to show that the ${\mathrm{O}}(n^{-1/2})$ term is ${\mathrm{o}}(n^{-1/2})$, with the exact rate depending on the covering number of $\mathcal{F}$}. This result shows that the term involving $\sigma_{\dagger,a}-\sigma_{\#,a}$ in (2) does not dominate in the Berry-Esseen-type bound in Theorem 3. If one is willing to ignore log factors when interpreting rates and ${E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}={\mathrm{O}}(n^{-1})$ (e.g., $\tilde{Q}_{a}$ is estimated in a parametric model), then $\sigma_{\dagger,a}-\sigma_{\#,a}={\mathrm{O}}(n^{-1/2})$ is at most comparable to $(\log n/n)^{1/2}$ in (2).
In many cases, the bounded uniform entropy integral condition $J(1,{\mathcal{F}},M)<\infty$ in Condition 10 is verified by establishing that $\mathcal{F}$} is subset of a VC or VC-hull class. I adopt the definition of VC index in van der Vaart and Wellner, (2023) and consider the following VC or VC-hull condition.
Condition 14 (VC or VC-hull).
@@
Under Part 1 of Condition 14, $\sup_{\mu}\log N(\epsilon M,{\mathcal{F}},L_{2}(\mu))\leq C_{1}(\nu)+2\nu\log(1/\epsilon)$ for all $\epsilon\in(0,1)$, where $C_{1}(\nu)$ is a constant depending on $\nu$ only (Theorem 2.6.7 in van der Vaart and Wellner,, 2023). Under Part 2 of Condition 14, $\sup_{\mu}\log N(\epsilon M,{\mathcal{F}},L_{2}(\mu))\leq C_{2}(\nu)\epsilon^{-2\nu/(2\nu+1)}$, where $C_{2}(\nu)$ is another constant depending on $\nu$ only (Corollary 2.6.12 in van der Vaart and Wellner,, 2023). These lead to the following result on the rates for (2), depending on which part of Condition 14 holds, and whether Conditions 11 and 12 hold. I will compare and interpret the rates in Section 3.3.
@@
Suppose that Conditions 15 and 10 hold. Consider the asymptotic scenario $n\to\infty$.